pt 1: How much aluminum is produced from 3350 kg of Al2O3?

pt2: The oxygen produced in Part 1 is allowed to react with carbon to produce carbon monoxide. Write a balanced equation describing the reaction of alumina with carbon. You need not include the states of matter in the balanced equation.

pt3: How much CO is produced from the alumina in Part 1?
Answer #1

In the problem, molar mass of Alumina=101.96g/mol

Molar mass of aluminum = 26.981539 amu

Molar mass of Carbon monoxide = 28.01g/mol

All these values are taken from wikipedia.
Answer #2

Part 1: Aluminum Production from Al₂O₃

Balanced Equation (Alumina Electrolysis):

2Al2O34Al+3O2

Given:

  • Mass of Al₂O₃ = 3350 kg

  • Molar mass of Al₂O₃ = 102 g/mol

  • Molar mass of Al = 27 g/mol


Steps:

  1. Moles of Al₂O₃:

    Moles=3350×103g102g/mol=32,843mol

  2. Moles of Al produced (2:4 ratio):

    Moles of Al=32,843×42=65,686mol

  3. Mass of Al:

    Mass=65,686mol×27g/mol=1,773,522g=1,774kg


Part 2: Balanced Equation (O₂ + C → CO)

2C+O22CO


Part 3: CO Production from Alumina

From Part 1:

  • Moles of O₂ produced (2:3 ratio):

    Moles of O2=32,843×32=49,265mol

From Part 2 (1:2 ratio for O₂:CO):

  1. Moles of CO:

    Moles of CO=49,265×2=98,530mol

  2. Mass of CO (molar mass = 28 g/mol):


  3. Mass=98,530mol×28g/mol=2,758,840g=2,759kg


Answers:

  1. Aluminum produced1,774 kg

  2. Balanced equation2C+O22CO

  3. CO produced2,759 kg


answered by: anonymous